Find the center of a circle with the equation x2 + y2 + 2x + 4y - 9 = 0
Answer by Nate(3500)

x^2 + y^2 + 2x + 4y - 9 = 0
x^2 + 2x + y^2 + 4y = 9
(x^2 + 2x) + (y^2 + 4y) = 9
(x + 1)^2 - 1 + (y + 2)^2 - 4 = 9
(x + 1)^2 + (y + 2)^2 = 14
C(-1,-2)